OpenStudy (anonymous):

Math Go over my answers only 5

OpenStudy (anonymous):

1.b 2.c 3.a 4.d 5.a Tell me if im wrong and please correct me @ehuman @SACAPUNTAS @Luigi0210

OpenStudy (anonymous):

recheck # 2, questions 3,4 and 5 will take more time to check than I have.

OpenStudy (anonymous):

@ehuman can you correct me?

OpenStudy (anonymous):

recheck # 2, questions 3,4 and 5 will take more time to check than I have.

OpenStudy (anonymous):

no like tell me what the correct answers? @ehuman

OpenStudy (anonymous):

for number 2, x=15.

OpenStudy (anonymous):

what about 1?

OpenStudy (anonymous):

you can have as many zeros as the degree of the function x = degree of 1 x^2 = degree of 2 x^3 = degree of 3 x^4 = degree of 4

OpenStudy (anonymous):

it is d?